Let W={(x,y,z)|x+y+z=0}

a, give a basis for W

b, what is the dimension of W


1
Expert's answer
2021-11-01T17:39:42-0400

Let "W=\{(x,y,z)|x+y+z=0\}."


a) Taking into account that


"W=\{(x,y,z)|x+y+z=0\}=\{(x,y,-x-y)|x,y\in\R\}\\=\{x(1,0,-1)+y(0,1-1)|x,y\in\R\},"


we conclude that the vectors "(1,0,-1)" and "(0,1,-1)" form a basis for "W."


b) Since two vectors form a basis for "W," the dimension of "W" is equal to 2.